Resource Supply and Usage, Special Order, Relevancy

Elliott, Inc., has four salaried clerks to process purchase orders. Each clerk is paid a salary of $24,200 and is capable of processing as many as 6,650 purchase orders per year. Each clerk uses a PC and laser printer in processing orders. Time available on each PC system is sufficient to process 6,650 orders per year. The cost of each PC system is $1,300 per year. In addition to the salaries, Elliott spends $28,196 for forms, postage, and other supplies (assuming 26,600 purchase orders are processed). During the year, 25,850 orders were processed.

Required:

1. Which of the resources associated with purchasing can be classified as flexible resources?

Forms, Postage and Supplies

Which of the resources associated with purchasing can be classified as committed resources?

PC system and clerks

2. Compute the total activity availability, and break this into activity usage and unused activity.

Activity availability: purchase orders
Activity usage: purchase orders
Unused activity: purchase orders

3. Calculate the total cost of resources supplied (activity cost), and break this into the cost of activity used and the cost of unused activity. When required, round your answers to the nearest dollar. If required, round your intermediate calculations to five decimal places.

Activity cost: $
Cost of activity used: $
Cost of unused activity:

$
Answer #1

solution forms, postage, and supplies Pc system and clerk. 26600 purchase orders Activity avastability Activity usage unusedWorking Notes :- Activity cost = 4(24200 + 1300)+(28196%25850/26600) = 129401 2) cost of unused activity = 4(24200 + 1300) *